OpenStudy (anonymous):

What is the simplified form of the following:

OpenStudy (anonymous):

\[\frac{ \frac{ x^2-1 }{ 9x } }{ \frac{ x^2+2x+1 }{ 3x^2 } }\]

OpenStudy (anonymous):

@hba @nincompoop

OpenStudy (anonymous):

I cant see equation right D:

OpenStudy (anonymous):

.

OpenStudy (anonymous):

I'll draw it for you... Hold on please! :)

OpenStudy (anonymous):



OpenStudy (anonymous):

Wow that is a hard one aint it XD @ranga

OpenStudy (ranga):

(x^2 - 1) = (x + 1)(x - 1) x^2 + 2x + 1 = (x + 1)^2 You can put those two factors and simplify then simplify the problem.

OpenStudy (anonymous):

I don't understand how to simplify that...

OpenStudy (jdoe0001):

you'd factor the numerators mainly

OpenStudy (jdoe0001):

and keep in mind that \(\bf \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\)

OpenStudy (ranga):

{ (x^2 - 1) / 9x } / { (x^2 + 2x + 1) / 3x^2 } = (x^2 - 1) / 9x * 3x^2 / (x^2 + 2x + 1) = (x + 1)(x - 1) / 9x * 3x^2 / (x + 1)^2 = (x + 1)(x - 1) / (x + 1)^2 * 3x^2 / 9x = You can try to simplify from here.

OpenStudy (anonymous):

so we'd have (x-1)/(x+1) * (3x^2)/9x?

OpenStudy (anonymous):

the second part would be x^2/3x after simplifying, right?

OpenStudy (ranga):

the second part you can cancel x too.

OpenStudy (anonymous):

so the final answer would be \[\frac{ x(x-1) }{ 3(x+1) }\]?

OpenStudy (anonymous):

Thank you sooooo much!!!

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

Do you think you could help me with a couple more if I post them?